Two bodies of masses 10kg and 4kg are moving in opposite directions with the same velocity of 5 meter per seconds .if they collide and stick together after the impact,calculate their velocity

1
Expert's answer
2022-03-08T08:56:03-0500

The law of conservation of momentum says

"m_1v_1+m_2v_2=(m_1+m_2)u"

The final velocity

"u=\frac{m_1v_1+m_2v_2}{m_1+m_2}"

"u=\frac{10*5+4*(-5)}{10+4}=2.14\:\rm m/s"